OpenStudy (nuccioreggie):

Simplify negative 3 over 4 divided by 2 over negative 8.

OpenStudy (whpalmer4):

Do you know how to divide by a fraction?

OpenStudy (whpalmer4):

\[\frac{\frac{-3}{4}}{\frac{2}{-8}}\]

OpenStudy (whpalmer4):

To divide by a fraction, flip the fraction over and multiply instead. \[\frac{\frac{a}{b} }{\frac{c}{d} }=\frac{a}{b}*\frac{d}{c}=\frac{a*d}{b*c}\]